# Are We Solving a Well-Defined Problem? A Task-Centric Perspective on Recommendation Tasks

### v2

@@ 6. Conclusion
While foundational works in recommender systems have been well established, there remains a lack of consensus within the community regarding baseline models and datasets. This gap is likely due to an insufficient understanding of RecSys tasks. In this paper, we provide an in-depth analysis of recommender system task formulations, emphasizing the need for clear and well-defined problem definitions to support effective evaluation and the development of more applicable solutions. We stress the importance of understanding input-output relationships in recommendation models, considering factors such as temporal dynamics and candidate item selection. Additionally, we examine the complexities of user-item interactions, including user-incurred costs during decision-making and the challenges posed by multi-step interactions in real-world scenarios. We also emphasize the goal of RecSys is to minimum (all forms of) user costs which may occur at any stage.
Many points discussed here may have been mentioned in previous literature or are common practices in the industry. Our goal, as a perspective paper, is not to provide a comprehensive survey. Instead, we aim to present a cohesive framework for better characterizing RecSys tasks, thereby guiding future academic research toward more meaningful contributions. This also calls for more details to be included in datasets released by the industry, as offline evaluations are particularly important for academic researchers who do not have access to A/B testing.
Ultimately, this paper seeks to clarify the relationship between task formulation, solution development, and evaluation (more specifically offline evaluation). By fostering a more structured and comprehensive understanding of these key elements, we hope to provide valuable insights for both new and experienced researchers, contributing to the continued advancement of the field of recommender systems across diverse real-world domains.

### v3

@@ 6. Conclusion
While foundational works in recommender systems are well established, there remains a lack of consensus within the community regarding baseline models and datasets. The reason behind is likely stemming from an insufficient understanding of RecSys task formulation. In this paper, we provide an in-depth analysis of recommendation tasks, emphasizing the need for clear and well-defined problem definitions to enable effective evaluation and the development of more practical solutions. We highlight the importance of understanding input-output relationships in recommendation models, accounting for factors such as temporal dynamics and candidate item selection. We further examine the complexities of user-item interactions, including user-incurred costs during decision-making and the challenges introduced by multi-step interactions in real-world settings. Ultimately, we argue that the central objective of RecSys is to minimize user costs across the entire recommendation process. The nature of these costs provides a basis for distinguishing among different recommendation tasks.
This paper aims to clarify the relationship between task formulation, solution design, and evaluation, particularly offline evaluation. By promoting a more structured and comprehensive understanding of these key elements, we hope to deepen the community’s appreciation of RecSys task complexity and support both new and experienced researchers in advancing the field across diverse real-world domains. We also urge researchers and reviewers to recognize the importance of task specificity and to value innovations tailored to distinct recommendation scenarios.